### v1

@@ Abstract
We look at the long-standing problem of segmenting unlabeled speech into word-like segments and clustering these into a lexicon. Several previous methods use a scoring model coupled with dynamic programming to find an optimal segmentation. Here we propose a much simpler strategy: we predict word boundaries using the dissimilarity between adjacent self-supervised features, then we cluster the predicted segments to construct a lexicon. For a fair comparison, we update the older ES-KMeans dynamic programming method with better features and boundary constraints. On the five-language ZeroSpeech benchmarks, our simple approach gives similar state-of-the-art results compared to the new ES-KMeans+ method, while being almost five times faster.
@@ V-A Comparison to Other Systems
The performance of all systems can be seen in Table I. When comparing the prominence-based method to ES-KMeans+, we see a tradeoff between NED and $R$-value: the simple method achieves better $R$-value on all languages except Wolof, while ES-KMeans+ gives better NED in all cases. Compared to the other approaches, both our systems achieve several state-of-the-art results and improve upon ES-KMeans. Prominence segmentation achieves a good tradeoff between metrics, finding a middle ground between ES-KMeans+ and DPDP (which consistently achieves a high $R$-value).
@@ V-B Effect of our Design Choices
TABLE III: Ablation study of the performance (%) of the main components of prominence segmentation with clustering on LibriSpeech dev-clean.
@@
MFCC + Prom. Seg. 75.3 65.3 29.7
HuBERT + SylSeg 41.9 52.2 16.3
HuBERT + Prom. Seg. 40.4 65.3 29.7

### v2

@@ Abstract
We look at the long-standing problem of segmenting unlabeled speech into word-like segments and clustering these into a lexicon. Several previous methods use a scoring model coupled with dynamic programming to find an optimal segmentation. Here we propose a much simpler strategy: we predict word boundaries using the dissimilarity between adjacent self-supervised features, then we cluster the predicted segments to construct a lexicon. For a fair comparison, we update the older ES-KMeans dynamic programming method with better features and boundary constraints. On the five-language ZeroSpeech benchmarks, our simple approach gives similar state-of-the-art results compared to the new ES-KMeans+ method, while being almost five times faster. Project webpage: https://s-malan.github.io/prom-seg-clus.
@@ V-A Comparison to Other Systems
The performance of all systems can be seen in Table I. When comparing the prominence-based method to ES-KMeans+, we see a tradeoff between NED and $R$-value: the simple method achieves better $R$-value on all languages except Wolof, while ES-KMeans+ gives better NED in all cases but English. Compared to the other approaches, both our systems achieve several state-of-the-art results and improve upon ES-KMeans. Prominence segmentation achieves a good tradeoff between metrics, finding a middle ground between ES-KMeans+ and DPDP (which consistently achieves a high $R$-value).
@@ V-B Effect of our Design Choices
TABLE III: Ablation scores (%) for the main components of prominence segmentation with clustering on LibriSpeech dev-clean. A boundary tolerance of 20 ms is allowed.
@@
MFCC + Prom. Seg. 75.3 50.7 15.6
HuBERT + SylSeg 41.8 39.1 7.5
HuBERT + Prom. Seg. 40.4 50.7 15.6